OpenStudy (anonymous):

square root n +10 = n +10

OpenStudy (anonymous):

\[\sqrt{n+10} = n+10\]

OpenStudy (saifoo.khan):

Take square on both sides. It will cancel out the radical. Leaving you with: \[n+10 = (n+10)^2\]

OpenStudy (saifoo.khan):

Multiplying will give you: n+10 = (n+10)(n+10) Using FOIL: n+10 = n^2 + 20n + 100 Can you solve now?